# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v1

@@ II Model
$H\!=\!\sum_{j_{z}}\left[h_{\text{2D}}\!+\!V(z)\sigma_{z}\right]c_{j_{z}}^{{\dagger}}c_{j_{z}},$ (1)
where $c_{j_{z}}^{{\dagger}}$ is the electron creation operator in the $j_{z}$-th layer, $h_{\text{2D}}\!=\!m_{z}\sigma_{z}\!+\!m_{0}[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)]\sigma_{z}\!+\!t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}$ represents the Chern insulator Hamiltonian when $m_{z}\times m_{0}<0$. Here, $a$ is the lattice constant, and $\sigma_{i}$ $(i\!=\!x,y,z)$ are Pauli matrices acting on the orbital space. The spatially periodic potential is given by $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $V_{z}$ is the potential strength, $z=j_{z}a$ is the position along the $z$-direction, and $\lambda_{z}\!=\!2\pi/Q_{z}$ is the periodic wavelength. In the following calculations, we use the fixed parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $\lambda_{z}\!=\!31a$ with $a\!=\!1$ nm. The detailed derivation of the matrix form for the real-space model in a periodic potential along the $z$-direction, with open boundary conditions (OBCs) in $y$-direction and periodic boundary conditions (PBCs) in $x$-direction, is provided in Section SI in the Supplemental Material Sup .
@@ II.1 Weyl-like model
$\!+i[{\cal A}_{k_{x}}^{mn}({\bf k},z),{\cal A}_{k_{y}}^{mn}({\bf k},z)],$
where ${\cal A}_{k_{\alpha}}^{mn}({\bf k},z)$ is the Berry connection of the occupied bands with $m,n\in{\rm occ}$, defined as
${\cal A}_{k_{\alpha}}^{mn}({\bf k},z)\!=\!i\langle\psi_{m}|\partial_{k_{\alpha}}\psi_{n}\rangle.$ (5)
@@ II.2 Weyl semimetal
$C(k_{z})\!=\!\frac{1}{4\pi}\int_{\text{B.Z.}}\!\!d^{2}{\bf k}\frac{\mathbf{h}({\bf k})}{|\mathbf{h}({\bf k})|^{3}}\cdot[\partial_{k_{x}}\mathbf{h}({\bf k})\times\partial_{k_{y}}\mathbf{h}({\bf k})],$ (9)
where ${\bf k}\!=\!(k_{x},k_{y})$ and the integral is over the 2D Brillouin zone. A detailed analytical expression for the Chern number (9) of the Weyl semimetal model (7) is provided in Section SIV in the Supplemental Material Sup .

### v2

@@ I Introduction
Off-resonant light serves as a powerful tool for inducing Floquet Weyl semimetals Wang et al. (2014); Hübener et al. (2017). Building on Floquet theory Oka and Aoki (2009); Zhan et al. (2024); Qin et al. (2023, 2022b, 2022a, 2024); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Dabiri et al. (2022); Cheraghchi and Askarpour (2023); Pervishko et al. (2018); Calvo et al. (2015); Seshadri and Sen (2022); Seshadri et al. (2019); Seshadri (2023); Lee et al. (2018); Magnus (1954); Blanes et al. (2009); Bukov et al. (2015); Eckardt and Anisimovas (2015); Chen et al. (2018c, d); Du et al. (2022); Wang et al. (2023), it would be intriguing to investigate the experimental implications of realistic periodic driving frequencies. This is particularly relevant in light of recent developments, such as the photoinduced anomalous Hall effect in graphene McIver et al. (2020) and its corresponding microscopic theory Sato et al. (2019), the experimental observation of Floquet states in graphene Merboldt et al. (2024), and the direct detection of Floquet-Bloch states in monolayer graphene Choi et al. (2024), WSe2 Aeschlimann et al. (2021), and the 3D topological insulator Bi2Se3 Wang et al. (2013). Additionally, experimental studies on Floquet transport effects, including anomalous linear transport in light-driven 3D Dirac electrons in bismuth Hirai et al. (2023) and nonperturbative nonlinear transport in a magnetic Floquet-Weyl semimetal Day et al. (2024), further highlight the potential of this approach.
@@ II Model
$H\!=\!\sum_{j_{z}}\left[h_{\text{2D}}({\bf k}_{||})\!+\!V(z)\sigma_{z}\right]c_{j_{z}}^{{\dagger}}c_{j_{z}},$ (1)
where $c_{j_{z}}^{{\dagger}}$ is the electron creation operator in the $j_{z}$-th layer, $h_{\text{2D}}({\bf k}_{||})\!=\!m_{z}\sigma_{z}\!+\!m_{0}[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)]\sigma_{z}\!+\!t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}$ represents the Chern insulator Hamiltonian when $m_{z}\times m_{0}<0$, and ${\bf k}_{||}\!=\!(k_{x},k_{y})$. Here, $a$ is the lattice constant, and $\sigma_{i}$ $(i\!=\!x,y,z)$ are Pauli matrices acting on the orbital space. The spatially periodic potential is given by $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $V_{z}$ is the potential strength, $z=j_{z}a$ is the position along the $z$-direction, and $\lambda_{z}\!=\!2\pi/Q_{z}$ is the periodic wavelength. In the following calculations, we use the fixed parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $\lambda_{z}\!=\!31a$ with $a\!=\!1$ nm. The detailed derivation of the matrix form for the real-space model in a periodic potential along the $z$-direction, with open boundary conditions (OBCs) in $y$-direction and periodic boundary conditions (PBCs) in $x$-direction, is provided in Section SI in the Supplemental Material Sup .
@@ II.1 Weyl-like model
$\!+i[{\cal A}_{k_{x}}^{mn}({\bf k}_{||},z),{\cal A}_{k_{y}}^{mn}({\bf k}_{||},z)],$
where ${\cal A}_{k_{\alpha}}^{mn}({\bf k}_{||},z)$ is the Berry connection of the occupied bands with $m,n\in{\rm occ}$, defined as
${\cal A}_{k_{\alpha}}^{mn}({\bf k}_{||},z)\!=\!i\langle\psi_{m}|\partial_{k_{\alpha}}\psi_{n}\rangle.$ (5)
@@ II.2 Weyl semimetal
$C(k_{z})\!=\!\frac{1}{4\pi}\int_{\text{B.Z.}}\!\!d^{2}{\bf k}_{||}\frac{\mathbf{h}({\bf k})}{|\mathbf{h}({\bf k})|^{3}}\!\cdot\![\partial_{k_{x}}\mathbf{h}({\bf k})\!\times\!\partial_{k_{y}}\mathbf{h}({\bf k})],$ (9)
where ${\bf k}_{||}\!=\!(k_{x},k_{y})$ and the integral is over the 2D Brillouin zone. A detailed analytical expression for the Chern number (9) of the Weyl semimetal model (7) is provided in Section SIV in the Supplemental Material Sup .